# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ A.1 Theorem 1:Analysis on subset GP:
The “Maximum Gradient Variance” principle for selecting the inducing points (or subsample set) aims to pick points that best capture the main gradient variations in the sample distribution, ensuring that $\mathbf{K}_{\mathcal{D}\mathcal{D}}\approx\hat{\mathbf{K}}$ accurately. As $M$ increases and the subset points are chosen more effectively, the approximation $\hat{\mathbf{K}}$ to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ improves, hence $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ decreases. Consider a set of $N$ samples $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and a corresponding kernel matrix: $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N},\quad(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j}),$ where $k$ is a positive definite kernel. Suppose we have observations $\mathbf{y}\in\mathbb{R}^{N}$ associated with these samples, and a probabilistic model (e.g., a GP) with parameters ${\theta}$ and the mean function $\mu$. The joint distribution of $\mathbf{y}$ given $\mathcal{D}$ and ${\theta}$ is: $\mathbf{y}\mid\mathcal{D},{\theta}\sim\mathcal{N}(\mu,\mathbf{K}_{\mathcal{D}\mathcal{D}}),$ where $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ is the covariance matrix induced by the kernel $k$. Define the gradient of the log-posterior (or log-likelihood) with respect to the latent function values $\mathbf{y}$: $g_{i}=\frac{\partial\log p(\mathbf{y}\mid\mathcal{D},{\theta})}{\partial y_{i}}=-(\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{y}-\mu))_{i}.$
@@
Now we force the initial batch of $n_{0}$ samples $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}$ and the latest sample $\mathbf{x}_{\mathrm{L}}$ to be included in $\mathcal{U}$. We then choose the remaining $M-n_{0}-1$ samples to maximize gradient variance. We select the $M-n_{0}-1$ samples that maximize the variance of gradient information. We have $\mathcal{U}\;=\;\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}\;\cup\;\{\mathbf{x}_{\mathrm{L}}\}\;\cup\;\mathcal{U}^{\prime},$ where $|\mathcal{U}^{\prime}|=M-n_{0}-1$ so that $|\mathcal{U}|=M$, the total subset $\mathcal{U}$ is still of size $M$. We consider a subset $\mathcal{U}\subset\mathcal{D}$ of size $M<N$ to build a low-rank approximation of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ are derived from $\mathcal{U}$. Let $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ be a positive definite kernel matrix with eigen-decomposition: $\mathbf{K}_{\mathcal{D}\mathcal{D}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ where $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{N}\geq 0$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N}]$. The best rank-$M$ approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ in spectral norm is: $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}=\mathbf{U}_{M}\bm{\Lambda}_{M}\mathbf{U}_{M}^{\top},$ where $\mathbf{U}_{M}=[\mathbf{u}_{1},\dots,\mathbf{u}_{M}]$ and $\bm{\Lambda}_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. By the Eckart–Young–Mirsky theorem: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|=\lambda_{M+1}.$ Suppose $\mathcal{U}\subset\mathcal{D}$, $|\mathcal{U}|=M$, produces a Nyström approximation: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$ If the subset $\mathcal{U}$ is chosen to approximate the principal eigenspace spanned by $\mathbf{U}_{M}$, then Nyström approximation theory (Drineas et al., 2005) guarantees that: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}=\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.$ where $\epsilon$ is an error control parameter related to the number of columns sampled randomly in the approximation.
@@
Let the UCB for the full GP model be defined as: $\mathrm{UCB}(\mathbf{x}_{*})=\mu(\mathbf{x}_{*})+\beta\sigma(\mathbf{x}_{*}),$ where $\mu(\mathbf{x}_{*})$ and $\sigma(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the full GP, respectively. For the sparse GP approximation, the UCB is given by: $\tilde{\mathrm{UCB}}(\mathbf{x}_{*})=\tilde{\mu}(\mathbf{x}_{*})+\beta\tilde{\sigma}(\mathbf{x}_{*}),$ where $\tilde{\mu}(\mathbf{x}_{*})$ and $\tilde{\sigma}(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the sparse GP approximation.
We aim to bound the error:
$|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|,$
in terms of the kernel matrix approximation error $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$. Given the bounds:
$|\Delta\mu(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$
$|\Delta\sigma(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|}.$
@@
$\bigl{|}\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})\bigr{|}\;=\;\bigl{|}\Delta\mu(\mathbf{x}_{*})+\beta\,\Delta\sigma(\mathbf{x}_{*})\bigr{|}\;\;\leq\;\;\bigl{|}\Delta\mu(\mathbf{x}_{*})\bigr{|}\;+\;\beta\,\bigl{|}\Delta\sigma(\mathbf{x}_{*})\bigr{|}.$
@@
$|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|+\beta\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|}.$
Using the Nyström approximation error bound $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq C^{\prime}\lambda_{M+1}$, the UCB error can be further bounded as:
$|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{1}(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})+\beta\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{1}(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})}.$

### v2

@@ A.1 Theorem 1:Analysis on subset GP:
The “Maximum Gradient Variance” principle for selecting the inducing points aims to pick points that best capture the main gradient variations in the sample distribution. As $M$ increases and the subset points are chosen more effectively, the approximation $\hat{\mathbf{K}}$ to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ improves, hence $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ decreases. Consider a set of $N$ samples $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and a corresponding kernel matrix: $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N},\quad(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j}),$ where $k$ is a positive definite kernel. Suppose we have observations $\mathbf{y}\in\mathbb{R}^{N}$ associated with these samples, and a probabilistic model (e.g., a GP) with parameters ${\theta}$ and the mean function $\mu$. The joint distribution of $\mathbf{y}$ given $\mathcal{D}$ and ${\theta}$ is: $\mathbf{y}\mid\mathcal{D},{\theta}\sim\mathcal{N}(\mu,\mathbf{K}_{\mathcal{D}\mathcal{D}}),$ where $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ is the covariance matrix induced by the kernel $k$. Define the gradient of the log-posterior (or log-likelihood) with respect to the latent function values $\mathbf{y}$: $g_{i}=\frac{\partial\log p(\mathbf{y}\mid\mathcal{D},{\theta})}{\partial y_{i}}=-(\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{y}-\mu))_{i}.$
@@
Now we force the initial batch of $n_{0}$ samples $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}$ and the latest sample $\mathbf{x}_{\mathrm{L}}$ to be included in $\mathcal{U}$. We then choose the remaining $M-n_{0}-1$ samples to maximize gradient variance. We select the $M-n_{0}-1$ samples that maximize the variance of gradient information. We have $\mathcal{U}\;=\;\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}\;\cup\;\{\mathbf{x}_{\mathrm{L}}\}\;\cup\;\mathcal{U}^{\prime},$ where $|\mathcal{U}^{\prime}|=M-n_{0}-1$ so that $|\mathcal{U}|=M$, the total subset $\mathcal{U}$ is still of size $M$. We consider a subset $\mathcal{U}\subset\mathcal{D}$ of size $M<N$ to build a low-rank approximation of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ are derived from $\mathcal{U}$. Let $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ be a positive definite kernel matrix with eigen-decomposition: $\mathbf{K}_{\mathcal{D}\mathcal{D}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ where $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{N}\geq 0$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N}]$. The best rank-$M$ approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ in spectral norm is: $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}=\mathbf{U}_{M}\bm{\Lambda}_{M}\mathbf{U}_{M}^{\top},$ where $\mathbf{U}_{M}=[\mathbf{u}_{1},\dots,\mathbf{u}_{M}]$ and $\bm{\Lambda}_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. By the Eckart–Young–Mirsky theorem: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|=\lambda_{M+1}.$ Suppose $\mathcal{U}\subset\mathcal{D}$, $|\mathcal{U}|=M$, produces a approximation: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
Gradient-based select $\mathcal{U}$. Here, we give a derivation showing that selecting the subset $\mathcal{U}$ by maximizing gradient-direction diversity is equivalent to the standard greedy Nyström column-selection. For a kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ and standard basis vectors $\bm{e}_{i}\in\mathbb{R}^{N}$, define $\bm{\phi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}\,\bm{e}_{i}\quad\text{(``kernel column" for sample $i$)},\qquad\bm{\psi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\bm{e}_{i}\quad\text{(``gradient column" for sample $i$)}.$ Suppose after $t$ steps we have selected index set $\mathcal{U}_{t}\subset\{1,\dots,N\}$. Let $\mathbf{C}_{t}=[\,\bm{\phi}_{j}\,]_{j\in\mathcal{U}_{t}}\in\mathbb{R}^{N\times t},\qquad\mathbf{G}_{t}=[\,\bm{\psi}_{j}\,]_{j\in\mathcal{U}_{t}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\mathbf{C}_{t}.$ Define the orthogonal projectors $\mathbf{P}_{t}^{(K)}=\mathbf{C}_{t}\,(\mathbf{C}_{t}^{\top}\mathbf{C}_{t})^{-1}\,\mathbf{C}_{t}^{\top},\qquad\mathbf{P}_{t}^{(g)}=\mathbf{G}_{t}\,(\mathbf{G}_{t}^{\top}\mathbf{G}_{t})^{-1}\,\mathbf{G}_{t}^{\top}.$ By construction, $\mathrm{range}(\mathbf{P}_{t}^{(K)})=\mathrm{range}(\mathbf{P}_{t}^{(g)})=\mathrm{span}\{\mathbf{C}_{t}\}.$ Using $\mathbf{G}_{t}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\mathbf{C}_{t}$ one checks $\mathbf{P}_{t}^{(g)}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\mathbf{C}_{t}\,\bigl{(}\mathbf{C}_{t}^{\top}\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-2}\mathbf{C}_{t}\bigr{)}^{-1}\,\mathbf{C}_{t}^{\top}\,\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1},$ and one can verify the key identity $\mathbf{I}-\mathbf{P}_{t}^{(g)}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\bigl{(}\mathbf{I}-\mathbf{P}_{t}^{(K)}\bigr{)}\,\mathbf{K}_{\mathcal{D}\mathcal{D}}.$ At iteration $t+1$: Gradient-diversity rule picks $i_{t+1}=\arg\max_{i\notin\mathcal{U}_{t}}\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(g)})\,\bm{\psi}_{i}\bigr{\rVert}.$ Nyström column-selection rule picks $i_{t+1}=\arg\max_{i\notin\mathcal{U}_{t}}\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\bigr{\rVert}.$ Using $\bm{\psi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\bm{e}_{i}$ and the conjugation relation, $\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(g)})\,\bm{\psi}_{i}\bigr{\rVert}=\bigl{\lVert}\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\bigr{\rVert}.$ Since $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}$ is a fixed invertible operator, maximizing $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\|$ is equivalent to maximizing $\|(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\|$. Therefore both rules select the same index $i_{t+1}$ at every step. Thus, selecting $\mathcal{U}$ by maximizing gradient-direction diversity coincides with the greedy Nyström column-selection.
If we choose $\mathcal{U}$ via a greedy procedure that maximizes gradient-information, then we defines the orthogonal projector onto the selected columns by $\mathbf{P_{\mathcal{U}}}\;=\;\mathbf{K_{D\,U}\,(K_{U\,U})^{-1}K_{U\,D}},$ and measures the subspace approximation error against the true principal eigenspace $\mathrm{span}(U_{M})$ as
$\begin{split}\mathbf{\epsilon_{g}}\;=\;\bigl{\|}\,(\mathbf{I}-\mathbf{P_{\mathcal{U}}})\,\mathbf{U_{M}}\bigr{\|}_{F}\;=\;\sqrt{\sum_{k=1}^{M}\bigl{\|}(\mathbf{I}-\mathbf{P_{\mathcal{U}}})\,\mathbf{u_{k}}\bigr{\|}^{2}}.\end{split}$ (10)
Then if $\mathcal{U}$ is chosen by Gradient Information, we obtain the analogous Nyström error bound:
$\begin{split}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\;\leq\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|\;+\;\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.\end{split}$ (11)
Random select $\mathcal{U}$. If the subset $\mathcal{U}$ is chosen to approximate the principal eigenspace spanned by randomly choose $\mathbf{U}_{M}$, then Nyström approximation theory [48] guarantees that: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}=\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.$ where $\epsilon$ is an error control parameter related to the number of columns sampled randomly in the approximation. In particular, if one samples $c$ columns uniformly at random , then for any $\delta\in(0,1)$ the following holds with probability at least $1-\delta$ [48]:
$\begin{split}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\;\leq\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|\;+\;\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2},\end{split}$ (12)
where $\epsilon\;=\;\sqrt{\frac{8}{c}\,\ln\frac{2M}{\delta}}\;=\;O\!\Bigl{(}\tfrac{1}{\sqrt{c}}\Bigr{)}.$
Comparison. Let $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ have eigenpairs $(\lambda_{k},u_{k})$, and let $U_{M}=[\,u_{1},\dots,u_{M}\,]\in\mathbb{R}^{N\times M}$ be the matrix of its top $M$ eigenvectors. Suppose we build a Nyström approximation by selecting columns greedily in $M$ steps; let $\mathcal{U}_{t}$ be the chosen indices after $t$ steps and $\mathbf{P_{t}}$ the corresponding orthogonal projector onto the span of those columns. Define the residual energy $E_{t}\;=\;\sum_{k=1}^{M}\|(\mathbf{I}-\mathbf{P_{t}})\mathbf{u_{k}}\|^{2}.$ Observe $\sum_{i=1}^{N}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}=\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{K}\|_{F}^{2}\;\geq\;\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\Lambda_{M}^{1/2}\|_{F}^{2}\;\geq\;\lambda_{1}\,\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\|_{F}^{2}=E_{t},$ where $\mathbf{\phi_{i}}=\mathbf{K}\,\mathbf{e_{i}}$ and $\Lambda_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. Hence among the remaining $N-t$ columns there is some index $j\notin\mathcal{U}_{t}$ with $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\geq\;\frac{1}{N-t}\sum_{i\notin\mathcal{U}_{t}}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}\;\geq\;\frac{E_{t}}{N-t}.$ Since the greedy rule picks precisely the column that maximizes $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|$, the new residual satisfies $E_{t+1}=E_{t}\;-\;\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\leq\;E_{t}\Bigl{(}1-\tfrac{1}{N-t}\Bigr{)}.$ Iterating the one-step bound from $t=0$ to $t=M-1$ gives $E_{M}\;\leq\;E_{0}\prod_{t=0}^{M-1}\Bigl{(}1-\tfrac{1}{N-t}\Bigr{)}=\;M\;\prod_{k=N-M+1}^{N}\Bigl{(}1-\tfrac{1}{k}\Bigr{)},$ since $E_{0}=\|U_{M}\|_{F}^{2}=M$. Using $1-\tfrac{1}{k}\leq e^{-1/k}$, we obtain $E_{M}\;\leq\;M\;\exp\!\Bigl{(}-\sum_{k=N-M+1}^{N}\tfrac{1}{k}\Bigr{)}\;\leq\;M\;\exp\!\bigl{(}-\tfrac{M}{N}\bigr{)}.$ Thus the final projection error $\epsilon_{g}=\sqrt{E_{M}}\;\leq\;\sqrt{M}\,\exp\!\bigl{(}-\tfrac{M}{2N}\bigr{)}=O\!\bigl{(}e^{-\tfrac{M}{2N}}\bigr{)}$. Because $\epsilon_{g}$ decays exponentially in $M$, it is asymptotically much smaller than the $O(1/\sqrt{M})$ decay of the random-sampling parameter $\epsilon$. Hence, greedy (gradient-based) selection yields a Nyström approximation error that is provably tighter than uniform random sampling for the same $M$.
The key insight is that each greedy step removes at least the average of the remaining projection error (or “energy”). If at step $t$ the total unexplained energy is $E_{t}$, then among the $N-t$ remaining columns there must be one whose residual contributes at least $E_{t}/(N-t)$. By selecting that column, the algorithm shrinks the leftover energy by a factor of $\bigl{(}1-\tfrac{1}{N-t}\bigr{)}$. Repeating this multiplicative reduction $M$ times drives the energy down by a product of $\bigl{(}1-\tfrac{1}{k}\bigr{)}$ terms, which is bounded by an exponential $e^{-M/N}$. Hence the projection error $\epsilon_{g}$ decays exponentially in $M$.
Our theoretical analysis dictates that the low-rank Nyström approximation error remains within a small relative tolerance of the full kernel. Formally, we require $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|=\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\leq 0.05\|\mathbf{K}_{\mathcal{D}\mathcal{D}}\|.$ Here, $\lambda_{M+1}$ denotes the $(M+1)$-th eigenvalue of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, and $\epsilon_{g}$ captures the projection error induced by the greedy column selection. We can solve for the smallest subset size $M_{\min}$ that satisfies this inequality.
@@
Let the UCB for the full GP model be defined as: $\mathrm{UCB}(\mathbf{x}_{*})=\mu(\mathbf{x}_{*})+\beta_{n}\sigma(\mathbf{x}_{*}),$ where $\mu(\mathbf{x}_{*})$ and $\sigma(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the full GP, respectively. For the sparse GP approximation, the UCB is given by: $\tilde{\mathrm{UCB}}(\mathbf{x}_{*})=\tilde{\mu}(\mathbf{x}_{*})+\beta_{n}\tilde{\sigma}(\mathbf{x}_{*}),$ where $\tilde{\mu}(\mathbf{x}_{*})$ and $\tilde{\sigma}(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the sparse GP approximation.
We aim to bound the error: $|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|,$ in terms of the kernel matrix approximation error $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$. Given the bounds:
$\begin{split}\bigl{|}\Delta\mu(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\|\mathbf{y}\|\,C_{M}\,\bigl{\|}\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\bigr{\|},\\ \bigl{|}\Delta\sigma(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\sqrt{C_{M}\,\bigl{\|}\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\bigr{\|}}.\end{split}$ (13)
@@
$\bigl{|}\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})\bigr{|}\;=\;\bigl{|}\Delta\mu(\mathbf{x}_{*})+\beta_{n}\,\Delta\sigma(\mathbf{x}_{*})\bigr{|}\;\;\leq\;\;\bigl{|}\Delta\mu(\mathbf{x}_{*})\bigr{|}\;+\;\beta_{n}\,\bigl{|}\Delta\sigma(\mathbf{x}_{*})\bigr{|}.$ (14)
@@
$|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|+\beta_{n}\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|}.$ (15)
Using the Nyström approximation error bound $\|\mathbf{K}-\hat{\mathbf{K}}\|\;\leq\;\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2},$ the UCB error can be further bounded as:
$\begin{split}\bigl{|}\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\|\mathbf{y}\|\,C_{M}\Bigl{(}\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr{)}\\ &\quad+\beta_{n}\|\mathbf{k}_{*\mathcal{D}}\|\,\sqrt{\,C_{M}\Bigl{(}\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr{)}}\,.\end{split}$ (16)